the wave numbers of specteral ilne is 5*10^5 m^-1 .the energy coressponding will be .

Grade:9

1 Answers

Mayajyothi C J
6 years ago
wave numer=\frac{1}{\lambda }
 
                     \frac{1}{\lambda }=\frac{1}{5\times 10^{5}}
                         ={0.2\times 10^{5}}
                 
                    E=\frac{hc}{\lambda }
                         =\frac{6.62\times 10^{-34}\times 3\times 10^{8}}{0.2\times 10^{-5} }
                       
                         =9.93\times 10^{-21}